OpenStudy (anonymous):

A ball at rest starts rolling down a hill with a constant acceleration of 3.2 meters/second2. What is the final velocity of the ball after 6.0 seconds? a) -9.6 meters/second b) 9.6 meters/second c) -19 meters/second d) 19 meters/second

OpenStudy (anonymous):

answer is d choice

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

\[apply v=u + at\] Here u=0 v=? a=3.2m/s^2 t= 6 sec